THE PURE ROTATIONAL SPECTRA OF ZnO IN THE EXCITED a^{3}Π_{i} STATE

L. N. ZACK, \underline{R. L. PULLIAM} and L. M. ZIURYS, Department of Chemistry, Department of Astronomy, Steward Observatory, University of Arizona, Tucson, AZ 85721.

The pure rotational spectra of ZnO in the excited a^{3}Π_{i} electronic state have been measured using direct absorption sub-millimeter techniques. This molecule was synthesized by reacting zinc vapor with N_{2}O in the presence of a DC discharge. Nine rotational transitions were recorded for the ^{64}ZnO, ^{66}ZnO, and ^{68}ZnO isotopomers in the v=0 state and data for the main isotopomer was measured in the v=1 state. All three spin components were observed for this state, each exhibiting lambda-doubling. The data were fit with a Hund case(a) Hamiltonian and rotational, spin-orbit, spin-spin, and lambda-doubling constants established. A bond length of 1.8436 Å was determined for this excited state, which is about 0.14 Å larger than that of the ground state.